# Better Late Than Never: Formulating and Benchmarking Recommendation Editing

### v1

@@ Abstract
1 Introduction
@@
Recommender systems (RSs) provide personalized recommendations by mining user preferences for items and have achieved great success in various applications Chen et al. (2023); Gao et al. (2023). Existing recommendation systems typically mine potential preferences from massive user historical behavior data , such as clicks and purchases, and then serve in online systems to recommend items. However, due to insufficient/unclean user behavior data Wang et al. (2021), limited model capabilities Zhao et al. (2024) as well as evolving user interests Zhang et al. (2020), recommendation systems may provide unsuitable recommendations to users and also reported by users. Among these unsuitable recommendation items, some may simply be uninteresting to users, while others can be more severe, causing user aversion or violating relevant laws and regulations. For example, recommending adult products to anonymous underage users or suggesting erroneous information to individuals with different ethnic or religious beliefs. Addressing this entails several requirements: (1) rapid rectification to mitigate negative impacts; (2) computation-efficient rectification to handle the potential frequent errors; (3) Preferably, solutions should work well even without training data access—a major plus for privacy—though access is allowed. Therefore, developing methods to quickly and efficiently rectify erroneous recommendations is crucial for fostering responsible and sustainable recommender systems in the real world.
Figure 1: An illustration of Recommendation Editing.
@@
To achieve recommendation rectification, we introduce recommendation editing task with a systematic framework, as shown in Figure 1. Given an explicit edit user-item pair, recommendation editing can edit the RS model with desired recommendation results. To fully understand and implement recommendation editing, it is essential to define three key goals: (1) Strict Rectification: unsuitable recommendation items, particularly those leading to serious issues like discrimination or illegal content, imperatively be removed after editing. Traditional methods like online/incremental recommendation Zhang and Kim (2023) or approaches considering negative feedback Huang et al. (2023) aim to iteratively refine and enhance the model with ongoing new data rather than strictly rectify known and unsuitable recommendation behavior without model retraining. Despite their infrequency, the significant negative impact of erroneous recommendations necessitates their prompt rectification. (2) Collaborative Rectification: the similar but unobserved erroneous recommendations items also imperatively be removed after editing. Given the significant negative impact of such recommendations, swift recommendation adjustments are necessary to correct these errors comprehensively and prevent similar future errors. While efficient fine-tuning strategies Wang et al. (2024) can address specific observed errors, they risk overfitting and may not generalize well to rectifying unobserved but related errors. (3) Concentrated Rectification: the majority of appropriate recommendations are preserved after editing, ensuring that unsuitable items from recommendations imperatively be removed only when necessary. Although addressing erroneous recommendations is important, large-scale, mature recommendation systems usually contain a small fraction of these severe errors. Retraining the model with recommendation error correction constraints can meet the above needs but might compromise the quality of the vast majority of acceptable recommendations. This compromise could detract from the overall user experience and jeopardize the system’s long-term viability.
@@
Although model editing is widely adopted in natural language processing (NLP) and computer vision (CV) Wang et al. (2023); Yao et al. (2023) to rectify model behaviors, it is a complex task in RS to rectify erroneous recommendations, and these solutions are not directly applicable to recommendation systems due to two key considerations. Firstly, the inherent nature of recommendation systems is non-iid data and involves a dynamic interaction between users and items, necessitating a sophisticated approach to loss functions that accurately captures this relationship. Secondly, the core objective of recommendation systems is to prioritize and rank items effectively, which requires the consideration of specialized ranking metrics. This distinction highlights the unique challenges in enhancing recommendation systems to avoid unsuitable suggestions, underlining the importance of a tailored approach for ‘recommendation editing’ that addresses these specific needs without the necessity for model retraining.
@@
To bridge the gap, this study concentrates on improving and evaluating the performance of recommendation editing. Our investigation unfolds across three critical dimensions. Firstly, we delve into the formulation of the recommendation editing problem and introduce three important editing objectives designed to evaluate the effectiveness of recommendation editing from various perspectives in Section 2.1. Subsequently, in Section 2.2, we explore the editing loss functions, a simple yet effective strong baseline, tailored for recommendation editing. Finally, a comprehensive benchmark with various baselines borrowing from other domains is thoroughly examined in Section 3 and  4. We demonstrate the effectiveness of our proposed methods after editing, thereby mitigating the issue of unsuitable recommendations. Our work aims to bridge the gap in current research by offering a comprehensive benchmark of RSs’ performance when tasked with the complex challenge of recommendation editing. We outline the major contributions below.
@@
  • We have established three key objectives for recommendation editing and have further developed three evaluation metrics to measure the performance of recommendation editing. These metrics enable a comprehensive assessment of its effectiveness.
@@
  • Our research benchmarks a series of recommendation editing methods, most of which draw inspiration from model editing techniques in the NLP field and incremental learning. Additionally, we propose a simple yet effective baseline method, named EFT, for recommendation editing and conduct a comparative analysis of various approaches.
2 Recommendation Editing
2.1 Problem Definition
@@
Recommendation Editing. Given the trained recommender system $f$ and the test data $\mathcal{D}_{test}$, the recommendation results can be denoted as =$\mathcal{R}=\{(u,i,\hat{y})_{r}\}$, where $(u,i,\hat{y})_{r}$ denotes the $r$-th recommendation results. However, some of the recommendation results may be inappropriate for certain users (such as different ages, races, or disabilities), which deserve to be edited to prevent negative impacts on users. We denote =$\mathcal{E}=\{(u,i,\tilde{y})_{e}\}\in\mathcal{R}$ as the set of user-item pairs that are expected to be edited, where $(u,i,\tilde{y})_{e}$ denotes the $e$-th editing pair. For same user/item pair $(u,i)_{r}=(u,i)_{e}$, the label is different $\hat{y}_{r}\neq\tilde{y}_{e}$.
@@
$\mathcal{E}^{\mathbb{E}}+\mathcal{E}^{\mathbb{I}}=\mathcal{E}\quad\text{and}\quad\mathcal{E}^{\mathbb{E}}+\mathcal{E}^{\mathbb{I}}+\bar{\mathcal{E}}=\mathcal{R}.$ (1)
Due to privacy protection and sparsity of user behavior, we often do not have access to the complete set of editing samples, instead, we only have access to a small explicit subset $\mathcal{E}^{\mathbb{E}}\in\mathcal{E}$. The recommendation editing aims to editing the recommendation output without retraining the entire model so that the following three properties can be satisfied:
@@
  • Prudent Editing: While recommendation editing is effective, the samples that need to be edited (explicitly or implicitly) typically represent only a small fraction of all the sample pairs. Otherwise, it would be preferable to retrain the model. Therefore, recommendation editing should be prudent and preserve recommendation results for $(u,i,\hat{y})_{i}\in\bar{\mathcal{E}}$.
2.2 Recommendation Editing Loss
@@
$x_{u,i}=P_{u}^{T}Q_{i},$ (2)
where $x_{u,i}$ is the score between user $u$ and item $i$. The final recommendation set $\mathcal{R}_{u}$ for user $u$ are selected as the items with top-$k$ highest scores among all the items. To achieve the goal of model-agnostic recommendation editing, we propose to directly operate on user embedding $\mathbf{P}$ and item embedding $\mathbf{Q}$, regardless of the way that learn these embeddings. It is worth noting that this loss function can also be directly used to optimize the whole recommendation model by backpropagating the gradients to the model parameters that obtain the representations.
In existing recommendation methods, Binary Cross-Entropy (BCE) and Bayesian Personalized Ranking (BPR) are the two most commonly used loss functions. The BCE loss is typically designed for the binary classification task. Traditional recommendation methods treat the user-item pair as a binary classification task where $y_{ui}=1$ denotes item $i$ is interesting for user $u$, and $y_{ui}=0$ otherwise. In the recommendation editing scenario, the editing pair provides the label $y_{ui}=0$ for the user-item pair. It is natural to optimize the corresponding user embedding $P_{u}$ and item embedding $Q_{i}$ by pushing them apart. However, in this way, we are only able to optimize and update the user embedding $P_{u}$ and item embedding $Q_{i}$, while the embedding of all the other users and items are not updated. As a result, only the item ranking for user $u$ and ranking of item $i$ in all other users will be updated, and the other user-item pairs: $\{(u^{\prime},i^{\prime}):u^{\prime}\neq u\ \text{and}\ i^{\prime}\neq i\}$ will not be updated. Concerning the comprehensive requirements of recommendation editing, although using the BCE loss may successfully edit the explicit editing pairs (high editing accuracy), the failure in editing implicit editing pairs $\mathcal{E}^{\mathbb{I}}$ makes it not suitable for recommendation editing.
The BPR loss Rendle et al. (2009b) is typically designed to optimize the relative ranking relationships between items for each user based on their preferences. In recommendation systems, we are more interested in the ranking of items by users rather than just binary classification predictions. As a result, we tend to utilize the idea of BPR loss to help edit the recommender system. In traditional recommendation, the explicit user feedback $(u,i)$ (such as user $u$ clicks or consumes an item $i$) is treated as positive item for the user $u$, while the negative pairs $(u,j)$ are usually sampled from implicit user feedback. The recommendation methods are usually optimized by maximizing the likelihood of correctly assigning higher score $x_{u,i}$ for the positive pair than other negative pairs $x_{u,j}$. In recommendation editing, given a few editing pairs $(u,i)$, it is natural to treat them as negative pairs. The core of utilizing BPR loss for editing lies in the definition of positive samples.
@@
$\mathcal{L}_{e\_bpr}=\sum_{(u_{e},i_{e})\in\mathcal{E}^{\mathbb{E}}}\sum_{j\in\mathcal{R}_{u_{e}}\setminus\{i_{e}\}}-log(\sigma(x_{u_{e},j}-x_{u_{e},i_{e}})).$ (4)
@@
3 The Setup of Recommendation Editing
Datasets and Editing Method. In our research, we utilize three datasets extensively employed in recommendation systems: Epinions [^1] https://cseweb.ucsd.edu/ jmcauley/datasets.html Zhao et al. (2014), KuaiRand [^2] https://kuairand.com/ Gao et al. (2022) and QB-video [^3] https://static.qblv.qq.com/qblv/h5/algo-frontend/tenrec_dataset.html Yuan et al. (2022) each featuring different types of negative feedback. The characteristics of these datasets are detailed in Table 1. To maintain dataset quality, we adhere to the commonly applied 10-core setting He et al. (2020) for preprocessing.
To assess the performance of different approaches, we have classified various methods into six categories relevant to incremental learning and model editing. Finetuning-based Approaches: These include FT (Fine-Tuning) Sinitsin et al. (2020) and EFT (Editing Fine-Tuning). Regularization-based Approaches: This category comprises LWF (Learning Without Forgetting) Li and Hoiem (2017), L2 (Online L2Reg) Lin et al. (2022), and SRIU (Sample Reweighting Incremental Update) Peng et al. (2021). Replay-based Approaches: Methods such as RSR (Random Sampling with a Reservoir) Vitter (1985) and SPMF (Stream-centered Probabilistic Matrix Factorization model) Wang et al. (2018) are included here. Optimization-based Approaches: SML (Sequential Meta-Learning) Zhang et al. (2020) represents this group. CF with Negative Feedback: This includes SiReN (Sign-aware Recommender system based on GNN models) Seo et al. (2022) and SiGRec (Signed Graph Neural Network Recommendation model) Huang et al. (2023). Editing Approach: EGNN (Editable Graph Neural Networks) Liu et al. (2023) and BiEGNN proposed by ourselves. To test the generalization of these editing methods, we conducted experiments across three representative recommendation models: MF (Matrix Factorization) Rendle et al. (2009a), LightGCN He et al. (2020) and XSimGCL (eXtremely Simple Graph Contrastive Learning) Yu et al. (2023). For further details on the methods and experimental settings, please refer to the Appendix A.
Editing Evaluation Metrics. Based on the characteristics of recommendation editing, we propose a comprehensive evaluation approach to assess the effectiveness. Specifically, we propose three metrics to evaluate the three proprieties above of recommendation editing:
    $EA=\frac{|\{r^{\mathbb{E}}_{ui}>k\}|}{|\mathcal{E}^{\mathbb{E}}|},\quad(u,i)\in\mathcal{E}^{\mathbb{E}}.$ (5)
    $EC=$ $\frac{1}{|\mathcal{E}^{\mathbb{I}}|}|\{r^{\mathbb{E}}_{ui}>k\ \&\ r_{ui}\leq k\}|+|\{r^{\mathbb{E}}_{ui}>r_{ui}\ \&\ r_{ui}>k\}|,\quad(u,i)\in\mathcal{E}^{\mathbb{I}}.$ (6)
    $EP=\frac{|\{r_{ui}<k\}\cap\{r^{\mathbb{E}}_{ui}<k\}|}{|\{r_{ui}<k\}\cup\{r^{\mathbb{E}}_{ui}<k\}|},\quad(u,i)\in\bar{\mathcal{E}}.$ (7)
To measure the overall erroneous editing performance, note that the number of implicit and unnecessary editing pairs ($\mathcal{E}^{\mathbb{I}}$ and $\bar{\mathcal{E}}$) could be imbalanced, we propose Editing Score (ES), the harmonic mean between EC and EP, to evaluate erroneous editing performance, i.e., $ES=2\frac{EC\cdot EP}{EC+EP}$  [^4] The harmonic mean in ES is similar to F1 penalizing large disparities between precision and recall, ensuring that both metrics need to be high to achieve a high score..
Table 1: Statistics of the experimental datasets.

### v2

@@ Abstract.
Recommender Systems, Collaborative Filtering, Model Editing
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
1. Introduction
@@
Recommender systems (RSs) provide personalized recommendations by mining user preferences for items and have achieved great success in various applications (Gao et al., 2023; Chen et al., 2023). Existing recommendation systems typically mine potential preferences from massive user historical behavior data, such as clicks and purchases, and then serve in online systems to recommend items. However, due to insufficient/unclean user behavior data (Wang et al., 2021), limited model capabilities (Fan et al., 2023) as well as evolving user interests (Zhang et al., 2020), recommendation systems may provide unsuitable recommendations to users and also reported by users. Among these unsuitable recommendation items, some may be uninteresting to users, while others can be more severe, causing user aversion or violating relevant laws and regulations. For example, recommending adult products to anonymous underage users or suggesting erroneous information to individuals with different ethnic or religious beliefs. Addressing this entails several requirements: (1) rapid rectification to mitigate negative impacts; (2) computation-efficient rectification to handle the potential frequent errors; (3) Preferably, solutions should work well even without training data access—a major plus for privacy—though access is allowed. Therefore, developing methods to quickly and efficiently rectify erroneous recommendations is crucial for fostering responsible and sustainable recommender systems in the real world.
Figure 1. Illustration of Recommendation Editing.
@@
To achieve recommendation rectification, we introduce recommendation editing task with a systematic framework, as shown in Figure 1. Given an explicit edit user-item pair, recommendation editing can edit RS model with desired recommendation results. To fully understand and implement recommendation editing, it is essential to define three key goals: (1) Strict Rectification: unsuitable recommendation items, particularly those leading to serious issues like discrimination or illegal content, imperatively be removed after editing. Traditional methods like online/incremental recommendation (Zhang and Kim, 2023) or approaches considering negative feedback (Huang et al., 2023) aim to iteratively refine and enhance the model with ongoing new data rather than strictly rectify known and unsuitable recommendation behavior without model retraining. Despite their infrequency, the significant negative impact of erroneous recommendations necessitates their prompt rectification. (2) Collaborative Rectification: the similar but unobserved erroneous recommendations items also imperatively be removed after editing. Given the significant negative impact of such recommendations, swift recommendation adjustments are necessary to correct these errors comprehensively and prevent similar future errors. While efficient fine-tuning strategies (Wang et al., 2023a) can address specific observed errors, they risk overfitting and may not generalize well to rectifying unobserved but related errors. (3) Concentrated Rectification: the majority of appropriate recommendations are preserved after editing, ensuring that unsuitable items from recommendations imperatively be removed only when necessary. Although addressing erroneous recommendations is important, large-scale, mature recommendation systems usually contain a small fraction of these severe errors. Retraining the model with recommendation error correction constraints can meet the above needs but might compromise the quality of the vast majority of acceptable recommendations. This compromise could detract from the overall user experience and jeopardize the system’s long-term viability.
@@
Although model editing is widely adopted in natural language processing (NLP) and computer vision (CV) (Yao et al., 2023; Wang et al., 2023b) to rectify model behaviors, it is a complex task in RS to rectify erroneous recommendations, and these solutions are not directly applicable to recommendation systems due to two key considerations. Firstly, the inherent nature of recommendation systems is non-iid(Non-Independent and Identically Distributed) data and involves a dynamic interaction between users and items, necessitating a sophisticated approach to loss functions that accurately captures this relationship. Secondly, the core objective of recommendation systems is to prioritize and rank items effectively, which requires the consideration of specialized ranking metrics. This distinction highlights the unique challenges in enhancing recommendation systems to avoid unsuitable suggestions, underlining the importance of a tailored approach for ‘recommendation editing’ that addresses these specific needs without the necessity for model retraining.
@@
To bridge the gap, this study concentrates on improving and evaluating the performance of recommendation editing. Our investigation unfolds across three critical dimensions. Firstly, we delve into the recommendation editing problem formulation and three evaluation metrics designed to gauge the effectiveness of recommendation editing from various perspectives in Section 3.2. Subsequently, in Section 4, we explore the editing loss functions, a simple yet effective strong baseline, tailored for recommendation editing. Finally, a comprehensive benchmark with various baselines borrowing from other domains is thoroughly examined in Section 5. We demonstrate the effectiveness of our proposed methods after editing, thereby mitigating the issue of unsuitable recommendations. Our work aims to bridge the gap in current research by offering a comprehensive benchmark of RSs’ performance when tasked with the complex challenge of recommendation editing. We outline the major contributions below.
@@
  • We have developed three evaluation metrics specifically designed to measure the performance of recommendation editing. These metrics enable a comprehensive assessment of its effectiveness.
@@
  • Our research benchmarks a series of recommendation editing methods, most of which draw inspiration from model editing techniques in the NLP field. Additionally, we propose a simple yet effective baseline method, named EFT, for recommendation editing and conduct a comparative analysis of various approaches.
2. Related Works
2.1. Online Recommendation
With the exponential growth of online information, recommender systems have become ubiquitous in various applications, helping users discover items aligned with their interests (Koren et al., 2021; He et al., 2017; Wang et al., 2019; He et al., 2020). In practice, RS operates in an incremental manner as new users, items, and user-item interactions are constantly being observed over time.
Online recommendation. To meet this demand, in recent years, some works have started focusing on efficient model updates for online learning in recommender systems. Among them, one kind of approach is based on continual learning, which typically requires maintaining a dynamic user model and making real-time adjustments based on user feedback and new data. ADER (Mi et al., 2020) periodically replays previous training samples to the current model with an adaptive distillation loss. Further, ReLoop (Cai et al., 2022) builds a self-correction learning loop for recommender systems with customized optimization loss based on user feedback. Another kind of approach is based on incremental learning, of which the goal is to quickly adapt current models to new data, and maintain high performance with minimal computational overhead through effective incremental learning algorithms or update strategies. MAN (Mi and Faltings, 2021) enhances a neural recommender system by incorporating a continuously queried and updated nonparametric memory. Then, FIRE (Xia et al., 2022) proposes a fast incremental learning method from the graph signal processing perspective.
Incremental recommendation. Existing online learning methods for recommendation rely on explicit user feedback. However, when containing incorrect items, it is preferable to proactively edit recommendations and ensure that incorrect items are removed before being recommended to the users, rather than expecting users to provide negative feedback after being exposed to such items.
2.2. Recommendation with Negative Feedback
The vast majority of recommendation systems make suggestions based on users’ historical positive feedback, neglecting users’ negative feedback. However, negative feedback is widely prevalent in real-world scenarios, such as low ratings, content skipping, and clicking the dislike button (Huang et al., 2023). It has been pointed out that utilizing negative feedback information can significantly enhance the performance of recommendation systems (Jeunen, 2019). To fully leverage the rich negative feedback, some methods have begun to explore how to combine positive and negative feedback for recommendations. For instance, movie recommendation systems (Chen et al., 2021) are proposed to establish a user positive profile and negative profile to enhance movie recommendations. Meanwhile, SiReN (Seo et al., 2022) and SiGRec (Huang et al., 2023) utilize both positive and negative feedback to construct a Signed Graph. By employing graph-based methods to integrate positive and negative feedback, these approaches aim to improve the performance of recommendation systems.
2.3. Model Editing
The concept of model editing was first introduced in the context of generative adversarial networks (Bau et al., 2020). Model edits were achieved by treating a neural network layer as a linear associative memory, allowing for the derivation of a low-rank update to the layer weights. This update was based on an input or a set of inputs, commonly referred to as a “concept”.
In addition, a variety of alternative editing algorithms have been proposed. For instance, (De Cao et al., 2021; Mitchell et al., 2022a) trained hyper-networks to update model weights. (Dai et al., 2022) identified “knowledge neurons” and then surgically modified them. (Mitchell et al., 2022b) augmented a target model with an auxiliary detector trained to identify inputs relevant to the editing task, and then diverted them to an edited model. See also (Hase et al., 2021) for an analysis of beliefs stored in language models and methods to update those beliefs. Recent work also extended model editing to open vocabulary image-language models (e.g., CLIP (Radford et al., 2021)), where it was demonstrated that interpolating between the original and edited weights allows users to navigate a trade-off between performance on the original and edited tasks (Ilharco et al., 2022).
The particular relevance to our work is the recent findings in (Lee et al., 2023): updating only a subset of model weights can result in greater robustness to distribution shift. We consider a different notion of robustness in this work, as well as a different set of models and editing/fine-tuning tasks than that work. We likewise update a more granular subset of weights ((Lee et al., 2023) updates “blocks” of weights, whereas we typically update weights in a single layer).
3. Recommendation Editing
In this section, we will begin by introducing the key notations used in this paper. Subsequently, we will formulate the recommendation editing problem and propose comprehensive evaluation metrics that consider the three properties of editing. Lastly, we will delve into a detailed discussion of the unique challenges posed by the recommendation editing problem.
3.1. Problem Definition
@@
Recommendation Editing. Given the trained recommender system $f$ and the test data $\mathcal{D}_{test}$, the recommendation results can be denoted as =$\mathcal{R}=\{(u,i,\hat{y})_{r}\}$, where $(u,i,\hat{y})_{r}$ denotes the $r$-th recommendation results. However, some of the recommendation results may be inappropriate for certain users (such as different ages, races, or disabilities), which deserve to be edited to prevent negative impacts on users. We denote =$\mathcal{E}=\{(u,i,\tilde{y})_{e}\}\subset\mathcal{R}$ as the set of user-item pairs that are expected to be edited, where $(u,i,\tilde{y})_{e}$ denotes the $e$-th editing pair. For same user/item pair $(u,i)_{r}=(u,i)_{e}$, the label is different $\hat{y}_{r}\neq\tilde{y}_{e}$.
@@
(1) $\mathcal{E}^{\mathbb{E}}+\mathcal{E}^{\mathbb{I}}=\mathcal{E}\quad\text{and}\quad\mathcal{E}^{\mathbb{E}}+\mathcal{E}^{\mathbb{I}}+\bar{\mathcal{E}}=\mathcal{R}.$
Due to privacy protection and sparsity of user behavior, we often do not have access to the complete set of editing samples, instead, we only have access to a small explicit subset $\mathcal{E}^{\mathbb{E}}\subset\mathcal{E}$. The recommendation editing aims to editing the recommendation output without retraining the entire model so that the following three properties can be satisfied:
@@
  • Prudent Editing: While recommendation editing is effective, the samples that need to be edited (explicitly or implicitly) typically represent only a small fraction of all the sample pairs. Otherwise, it would be preferable to retrain the model. Therefore, model editing should be prudent and preserve recommendation results for $(u,i,\hat{y})_{i}\in\bar{\mathcal{E}}$.
Note that in practice, the recommender system is usually trained on large-scale user behavior data and has a large number of parameters, we can not re-train or fine-tune the model since we do not have access to the model parameters as well as training data.
3.2. Editing Evaluation Metrics
Based on the above properties, we propose a comprehensive evaluation approach to assess the effectiveness of the recommendation editing. Specifically, we propose three metrics to evaluate the three aforementioned proprieties of recommendation editing:
    (2) $EA=\frac{|\{r^{\mathbb{E}}_{ui}>k\}|}{|\mathcal{E}^{\mathbb{E}}|},\quad(u,i)\in\mathcal{E}^{\mathbb{E}}.$
    (3) $EC=$ $\frac{1}{|\mathcal{E}^{\mathbb{I}}|}|\{r^{\mathbb{E}}_{ui}>k\ \&\ r_{ui}\leq k\}|$
    +|$+|\{r^{\mathbb{E}}_{ui}>r_{ui}\ \&\ r_{ui}>k\}|,\quad(u,i)\in\mathcal{E}^{\mathbb{I}}.$
    (4) $EP=\frac{|\{r_{ui}<k\}\cap\{r^{\mathbb{E}}_{ui}<k\}|}{|\{r_{ui}<k\}\cup\{r^{\mathbb{E}}_{ui}<k\}|},\quad(u,i)\in\bar{\mathcal{E}}.$
To measure the overall erroneous editing performance, note that the number of implicit and unnecessary editing pairs ($\mathcal{E}^{\mathbb{I}}$ and $\bar{\mathcal{E}}$) could be imbalanced, we propose Editing Score (ES), the harmonic mean between EC and EP, to evaluate erroneous editing performance, i.e. $ES=2\frac{EC\cdot EP}{EC+EP}$  [^1] The harmonic mean in ES is similar to F1 penalizing large disparities between precision and recall, ensuring that both metrics need to be high to achieve a high score..
3.3. Discussions
Given the definition and the three key goals of recommendation editing, in this subsection, we discuss the relations and differences between recommendation editing with many existing methods.
Online/Incremental Recommendation aims to continuously update the recommendations model based on the new users’ feedback. While online/incremental recommendation methods can be adapted to handling negative feedback, the goals are quite different from model editing. Specifically, online/incremental recommendation focuses on model adaption to new user feedback for accurate recommendations for all users rather than editing those important but erroneous recommendation items. Consequently, the erroneous recommendation items that require editing may be compromised using online/incremental recommendation methods due to the preference modeling of a massive number of users. Model Editing sets strictly rectifying erroneous recommendation items as the primary goal, while online aims at adapting the new user feedback.
Recommendation with Negative Feedback focuses on modeling user preferences by considering both positive and negative feedback. However, its effectiveness relies on training from scratch and thus requires access to all historical feedback data. This requirement is impractical in recommendation editing contexts, due to the inherent inefficiencies and the common issue of incomplete datasets.
Recommendation unlearning seeks to eliminate the influence of specific training data from the model, whereas recommendation editing aims to avert erroneous recommendations generated by the recommendation model. The divergent goals of these approaches pose challenges in applying recommendation unlearning strategies to meet the strict rectification goal of recommendation editing.
Model Editing has garnered attention in computer vision and natural language processing. However, it overlooks the inherent nature of recommendation systems involving non-independent and identically distributed (non-iid) data and dynamic interactions between users and items. Furthermore, model editing primarily targets instance-wise tasks, such as classification. In recommendation editing, the ranking task involves the relative importance of items for each user. This distinction emphasizes the need for recommendation editing approaches specifically designed to address the nuanced challenges of ranking.
4. Recommendation Editing Loss
@@
(5) $x_{u,i}=P_{u}^{T}Q_{i},$
where $x_{u,i}$ is the score between user $u$ and item $i$. The final recommendation set $\mathcal{R}_{u}$ for user $u$ is selected as the items with top-$k$ highest scores among all the items. To achieve the goal of model-agnostic recommendation editing, we propose to directly operate on user embedding $\mathbf{P}$ and item embedding $\mathbf{Q}$, regardless of the way that learn these embeddings. It is worth noting that this loss function can also be directly used to optimize the whole recommendation model by backpropagating the gradients to the model parameters that obtain the representations.
In existing recommendation methods, Binary Cross-Entropy (BCE) and Bayesian Personalized Ranking (BPR) are the two most commonly used loss functions. The BCE loss is typically designed for the binary classification task. Traditional recommendation methods treat the user-item pair as a binary classification task where $y_{ui}=1$ denotes item $i$ is interesting for user $u$, and $y_{ui}=0$ otherwise. In the recommendation editing scenario, the editing pair provides the label $y_{ui}=0$ for the user-item pair. It is natural to optimize the corresponding user embedding $P_{u}$ and item embedding $Q_{i}$ by pushing them apart. However, in this way, we are only able to optimize and update the user embedding $P_{u}$ and item embedding $Q_{i}$, while the embedding of all the other users and items is not updated. As a result, only the item ranking for user $u$ and ranking of item $i$ in all other users will be updated, and the other user-item pairs: $\{(u^{\prime},i^{\prime}):u^{\prime}\neq u\ \text{and}\ i^{\prime}\neq i\}$ will not be updated. Concerning the comprehensive requirements of recommendation editing, although using the BCE loss may successfully edit the explicit editing pairs (high editing accuracy), the failure in editing implicit editing pairs $\mathcal{E}^{\mathbb{I}}$ makes it not suitable for recommendation editing.
The BPR loss (Rendle et al., 2012) is typically designed to optimize the relative ranking relationships between items for each user based on their preferences. In recommendation systems, we are more interested in the ranking of items by users rather than just binary classification predictions. As a result, we tend to utilize the idea of BPR loss to help edit the recommender system. In traditional recommendation, the explicit user feedback $(u,i)$ (such as user $u$ clicks or consumes an item $i$) is treated as a positive item for the user $u$, while the negative pairs $(u,j)$ are usually sampled from implicit user feedback. The recommendation methods are usually optimized by maximizing the likelihood of correctly assigning a higher score $x_{u,i}$ for the positive pair than other negative pairs $x_{u,j}$. In recommendation editing, given a few editing pairs $(u,i)$, it is natural to treat them as negative pairs. The core of utilizing BPR loss for editing lies in the definition of positive samples.
@@
(7) $\mathcal{L}=\sum_{(u_{e},i_{e})\in\mathcal{E}^{\mathbb{E}}}\sum_{j\in\mathcal{R}_{u_{e}}\setminus\{i_{e}\}}-log(\sigma(x_{u_{e},j}-x_{u_{e},i_{e}})).$
@@
5. Experiments
In this section, we conduct comprehensive experiments to answer the following research questions:
  • RQ1:
    How effective are the various editing methods in terms of three editing goals?
  • RQ2:
    How quickly can editing methods rectify recommendation errors?
  • RQ3:
    What is the impact of the choice of editing pairs on the performance of editing methods?
  • RQ4:
    How should an appropriate objective for editing methods be determined?
5.1. Experimental Setup
5.1.1. Datasets
In our study, we extensively utilize three datasets frequently employed in research on recommendation systems, each distinguished by varying types of negative feedback. This approach is adopted because the intensity of negative feedback in real-world applications tends to be unpredictable and must be tailored according to specific business needs.
Epinions [^2] https://cseweb.ucsd.edu/ jmcauley/datasets.html (Zhao et al., 2014), sourced from online consumer review platform Epinions.com, comprises user ratings with a range from 1 to 5, detailed item reviews, and directed trust relationships among users. We classify interactions as positive feedback if ratings exceed 3, and negative feedback if ratings are 3 or lower.
KuaiRand [^3] https://kuairand.com/ (Gao et al., 2022), an unbiased sequential recommendation dataset, is derived from the recommendation logs of Kuaishou, a popular video-sharing mobile application. A user clicking on a video indicates positive feedback, while negative feedback is represented by items shown to a user who then clicks the dislike button.
QB-video [^4] https://static.qblv.qq.com/qblv/h5/algo-frontend/tenrec_dataset.html (Yuan et al., 2022) is originated from the QQ browser. The positive feedback here is represented by the user clicking on videos, and negative feedback occurs when a user is shown a video but does not click on it.
The statistics of these datasets are shown in Table 1. To ensure the quality of the datasets, we follow the widely used 10-core setting (Wang et al., 2019) for dataset preprocessing.
Table 1. Statistics of the experimental datasets.